OpenStudy (anonymous):

Find the slope of a line parallel to the following line: 5y + 3x =1

OpenStudy (anonymous):

slope of \[ax+by=c\] is \[-\frac{a}{b}\]

OpenStudy (anonymous):

with the formula, or simply turn into the regular form y=mx+q, so y=-3/5x+1/5 so the slope is -3/5

OpenStudy (anonymous):

THANK YOU SO MUCH :D

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

its y=mx + B